10        20        30 1ba1 T0089  ----MSKGPAVGIDLGTTYSKVGVFQ--HGKVEIIANDQgnrttpsyvaftdterligdaaknqvamnpt MIDLSKTVFYTSIDIGSRYIKGLVLGKRDQEWEALAFSS............................... 40        50        60        70        80        90       100 1ba1 T0089  ntvFDAKRLIGRRFDD--------------AVVQSDMKHWPFMVVNDAGR------PKVQVEYKGETKSF ...VKSRGLDEGEIKDAIAFKESVNTLLKELEEQLQKSLRSDFVISFSSVSFEREDTVIERDFGEEKRSI 110       120           130       140       150       160       170 1ba1 T0089  YPEEVSSMVLTKMKEiaeaYLGKTVTNAVVTVPAYFNDSQR---------------------------QA TLDILSEMQSEALEK....LKENGKTPLHIFSKRYLLDDERIVFNPLDMKASKIAIEYTSIVVPLKVYEM 180       190       200         210       220       230 1ba1 T0089  TKDAGTIAGLNVLRIINEPTAAAIAYGLDKkvGAERNVLIFDLGGGTFDVSILTIEDGIFevkstAGDTH FYNFLQDTVKSPFQLKSSLVSTAEGVLTTP..EKDRGVVVVNLGYNFTGLIAYKNGVPIK.....ISYVP 240                          250       260       270       280 1ba1 T0089  LGGEDFDNRMVnhfiaefkrkhkkdisenkRAVRRLRTACERAKRTLSSST---QASIEIDSLYEGIDFY VGMKHVIKDVS...................AVLDTSFEESERLIITHGNAVYNDLKEEEIQYRGLDGNTI 290       300       310       320       330       340       350 1ba1 T0089  TSITRARFEELNADLFRGTLDPVEKALRDAKL------DKSQIHDIVLVGGSTRIPKIQKLLQDFFnGKE KTTTAKKLSVIIHARLREIMSKSKKFFREVEAKIVEEGEIGIPGGVVLTGGGAKIPRINELATEVF.KSP 360       370       380       390       400       410 1ba1 T0089  LNKS-----------------INPDEAVAYGAAVQAAILSGDKSE------------------- VRTGCYANSDRPSIINADEVANDPSFAAAFGNVFAVSENPYEETPVKSENPLKKIFRLFKELME

Figure 1: Comparison of candidate ../1ba1/1ba1-T0089-global.pw.a2m (lines) to reference ../1ba1/T0089-1ba1-2track-global.pw.a2m (letters). Reference alignment has 321 matching columns, Candidate alignment has 307 matching columns. 32 were used (for template shift+target shift) to compute shift score. The shift score is 0.005988, with a min shift of 167, max shift of 0, and mean of 0.000000. Optimal subalignment has 166 matching columns, with 238 paired matches. 1 were used (for template shift+target shift) to compute shift score. Optimum shift score is 0.542915, with a min shift of 545, max shift of 0, and mean of 2.895779. Optimum achieved by removing the 212 columns indicated with dotted lines from the candidate alignment.